OpenStudy (anonymous):

1. -6+5(-1-b)=19 2. -14=-(-2x+2) 3.-9+4r=4r-3-6 Please explain it all :]

OpenStudy (saifoo.khan):

1. \[-6 + 5(-1-b) = 19\]\[5(-1-b) = 19 + 6\]\[5(-1-b) = 25\]\[-1-b = 25/5\]-1-b = 5

OpenStudy (saifoo.khan):

-b = 5 + 1 -b = 6 b = -6